OpenStudy (desmarie):

A circle in the xy‑plane has the equation (x+92)^2+(y+54)^2=73. What is the length of the diameter of the circle?

OpenStudy (welshfella):

the general form of a circle is (x - a)^2 + (y - b)^2 = r^2 where r = radius

OpenStudy (desmarie):

thanks :)

OpenStudy (welshfella):

yw

OpenStudy (desmarie):

since they already included the 73 am I suppose to still square that or is that the official answer?

OpenStudy (desmarie):

@welshfella @jim_thompson5910

OpenStudy (rayaz):

The 73 has been obtained by squaring sqrt(73) so there is no need to square it again.

OpenStudy (desmarie):

ok thanks ....sorry that I cant medal you I already chose someone else

OpenStudy (rayaz):

No problem. Did you understand how to solve it?

OpenStudy (desmarie):

ehhh

jimthompson5910 (jim_thompson5910):

match up the r^2 and the 73 this means \[\Large r^2 = 73\] \[\Large r = \sqrt{73}\]

jimthompson5910 (jim_thompson5910):

That will give you the radius. Double it to get the diameter.

OpenStudy (desmarie):

geez jim.....you are a GENIUS!!

OpenStudy (desmarie):

so the answer is 2rad73......correct?